Consider a ball of mass m1 moving with a velocity u1 collides elastically with a ball of mass m2 at rest.Show that the maximum energy transfer takes place when m1=m2.
1
Expert's answer
2011-09-06T07:02:04-0400
The Energy transfer equation is
W=E2 - E1=m (u12)/2 - 0 = m (u12)/2.


The Momentum conservation law
m1u1 = m2u2+m1 u1'.


Thus, when masses are equal, the second ball begins to move with the speed u2=u1 - u1' .